OpenStudy (anonymous):

Graphing quadratics by factoring; Quadratics and graphing are my two worst things, could someone try to explain to me step by step how to do this? Here is an example: x^2=14x–49

OpenStudy (saifoo.khan):

\[x^2 -14x = -49\]\[x^2 - 14x + (\frac{14}{2})^2 - (\frac{14}{2})^2 = -49\]\[(x-7)^2-(7)^2 = -49\] \[(x-7)^2 = 0\]

OpenStudy (saifoo.khan):

So, Turning point = (7,0)